Circular No. 8057

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003P IN MCG +09-13-107 Further to IAUC 8030, T. Boles reports the discovery of an appar-

ent supernova (mag 18.0) on unfiltered CCD images taken on Jan. 23.916, 24.250, and 25.758 UT. The new object is located at α = 8^{h}01^{m}15^{s}.63, δ = +55^{o}44′34′′.8, which is ≈ 3′′.5 east and 3′′.5 north of the center of MCG +09-13-107. SN 2003P is not present on Boles’ images from 2002 Nov. 3 (limiting mag 19.5) or Apr. 8 (limiting mag 19.0), and it is not present on second Palomar Sky Survey red (1995) or blue (1998) plates or on the Quick V northern plate (1984).

SUPERNOVAE 2003L, 2003M, AND 2003O S. Valenti and E. Cappellaro, Universitá di Napoli; S. Danese and

G. Di Pede, Universitá di Padova; and H. Navasardyan, A. Pastorello, S. Benetti, and M. Turatto, Istituto Nazionale di Astrofisica, Padova, report that inspection of a reduced spectrum (range 365–770 nm; resolution 2.3 nm), taken at the Asiago 1.82-m telescope (+ AFOSC) on Jan. 25.00 UT, shows that SN 2003L (cf. IAUC 8048) is of type Ic, a few days after maximum. The spectrum shows a relatively blue continuum and is dominated by strong P-Cyg lines of Ca II (H and K) and Fe II. A relatively weaker Si II 635.5-nm line is also visible. A low expansion velocity is derived from the Si II line (∼ 5900 km/s). The redshift of the galaxy at the supernova location, derived from a strong and narrow Hα emission superimposed on the supernova spectrum, is ∼ 6720 km/s.

Benetti et al. also obtained a spectrum (as above; range 365–900 nm; resolution 3.5 nm at wavelength, λ, > 770 nm) of SN 2003M: “The spectrum, after correction for the redshift of the parent galaxy, reveals a strong resemblance to that of SN 1994I around maximum light for λ > 500 nm, but it shows weak Fe lines at λ < 500 nm. This confirms the physical association of the supernova with the elliptical galaxy UGC 7224, even if the supernova is located ∼ 30000 pc from its nucleus. This is possibly the first detection of a core-collapse supernova in an early-type galaxy.”

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003O (cf. IAUC 8055), obtained by M. Calkins on Jan. 25.12 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-II supernova. Adopting the NASA/IPAC Extragalactic Database recession velocity of 4941 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 9100 km/s.

2003 January 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green